OpenStudy (anonymous):

Use the Binomial Theorem to find the coefficient of the x^2 y^6 term in the expansion of (5x-2y)^8.

jimthompson5910 (jim_thompson5910):

Hint: Compute 8 C 6

jimthompson5910 (jim_thompson5910):

Then multiply this by (5^2)*(2^6) = 1600

OpenStudy (anonymous):

?????

jimthompson5910 (jim_thompson5910):

Do you know how to evaluate 8 C 6?

OpenStudy (anonymous):

no >.>

jimthompson5910 (jim_thompson5910):

n C r = (n!)/(r!(n-r)!) 8 C 6 = (8!)/(6!*(8-6)!) 8 C 6 = (8!)/(6!*2!) 8 C 6 = (8*7*6!)/(6!*2!) 8 C 6 = (8*7)/(2!) 8 C 6 = (8*7)/(2*1) 8 C 6 = (56)/(2) 8 C 6 = 28

jimthompson5910 (jim_thompson5910):

Multiply that result with 1600 to get 28*1600 = 44800 So the coefficient of the x^2y^6 term is 44800

OpenStudy (anonymous):

oh